# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

### v2

@@ 4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class (12 categories of the taxonomy +“Other”) classification of WildFireCan-MMD. In general, we employ VLMs as zero-shot classifiers, and compare the results with fine-tuned transformer-based classifiers as well as to classical methods such as decision tree, and Gaussian naive bayes, as minimum baselines.
Since each post on X can have up to 4 associated images, we duplicated the text and paired with each image as an individual sample. For training and testing the custom classifiers, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. We use the same test set for all of the classifiers, including zero-shot VLMs.
4.1 VLMs as Zero-Shot Classifiers:
We considered three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025) and tested them in a zero-shot setting. We created classification prompts for all three VLMs where, with each query to the model, a list of all categories and their descriptions is given as a prompt along with a data instance (text + image) to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
gpt-4o-mini: We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text and image from the tweet, and user prompt. The responses are saved as a file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allows for the calculation of metrics.
LLaVA: We opted for the 13B 4-bit version of LLaVA due to hardware constraints. Using a combination of a publicly available Colab notebook  [^4] https://github.com/camenduru/LLaVA-colab and the original code from the open-source LLaVA repository  [^5] https://github.com/haotian-liu/LLaVA, the model was loaded onto the GPU within Colab. The answers are saved similarly as described for gpt-4o-mini.
smolVLM: Being the smallest of the VLMs experimented with, only requiring <10GB of GPU ram, experiments were able to be performed locally without quantization. Due to its inclusion in the Transformers library, loading and querying the model with the previously created prompts was simpler and cheaper than the other methods, while taking a similar amount of time to process.
4.2 Custom Classifiers
Prior work by Koshy and Elango (2023) demonstrated that a multimodal architecture combining RoBERTa with a BiLSTM and ViT, using multiplicative fusion and an additional attention mechanism, achieved the best performance on the CrisisMMD dataset. We evaluated this architecture, among several others, on our WildFireCA-MMD dataset. Our findings indicate that the BiLSTM and additional attention components did not improve performance in our context. Instead, a simpler architecture employing early fusion via feature concatenation of RoBERTa and ViT embeddings yielded the best results. We experimented with several model structures described in the following.
2-head structure:This model consists of two parallel architectures, an image head and a text head. The image head uses the ViT666’google/vit-base-patch16-384’ transformer, and the text head uses the RoBERTa777’roberta-base’ transformer. A linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion.
3-head structure: To enhance classification without more manual labeling, we extracted text from images using EasyOCR [^8] https://github.com/JaidedAI/EasyOCR. The extracted text was saved in a new column (extracted_text), defaulting to a filler if none was found. Our classifier was updated to include a third RoBERTa head for this column, as shown in Figure D.2.
Training and Evaluation: A random seed was used for reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by our ViT model). We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (~1000 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
4.3 Baseline Classifiers
To establish baseline, we evaluated several classic classifiers on our dataset, namely Decision Tree, Gaussian Naive Bayes, and K-Nearest Neighbors (KNN). For textual data, we extracted TF-IDF feature matrices by applying the TfidfVectorizer with max_features=5000. The image data, was resized to 64$\times$64 pixels. The resulting TF-IDF and image matrices were concatenated and used as input to the classifiers. Hyperparameters were set to default values in Scikit-learn  [^9] https://scikit-learn.org/stable/.
@@ 5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and baseline models per class and as an average over all classes. Also, Figures E.1 and E.2 show the confusion matrices of the custom classifiers and VLMs, respectively. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Some classes are harder to distinguish than others. In general, we observe a large variance of classification F-scores across classes. For example, all classifiers struggle with classifying examples of Evacuees and General Information. We explain this observation by the great variability of information in these classes. On the other hand, classes with less information variety, such as Insurance and Smoke and Air Quality, are more manageable for most classifiers to detect.
Table 2: F1-scores per-class for VLMs, custom and baseline classifiers. Maximum results are in bold.
VLMs Custom Classifiers Baseline Classifiers
Class LLaVA GPT-4o-mini smolVLM 2-head 3-head Dec. Tree Gauss. NB KNN
Evacuees 0.17 0.37 0.25 0.77 0.73 0.47 0.53 0.29
General Information 0.14 0.16 0.00 0.59 0.46 0.23 0.46 0.21
Preparedness 0.61 0.65 0.23 0.82 0.81 0.57 0.72 0.49
Weather Reports 0.59 0.64 0.00 0.88 0.91 0.46 0.63 0.52
Warnings & Status Updates 0.36 0.45 0.00 0.80 0.76 0.45 0.59 0.43
Reports of Actions of Responders 0.29 0.63 0.00 0.89 0.86 0.45 0.70 0.40
Infrastructure 0.38 0.40 0.00 0.92 0.95 0.47 0.75 0.52
Political 0.70 0.77 0.00 0.84 0.84 0.52 0.75 0.31
Insurance 0.93 0.90 0.92 0.93 0.89 0.93 0.93 0.89
Advertisement 0.08 0.68 0.00 0.76 0.76 0.22 0.65 0.40
Smoke and Air Quality 0.76 0.79 0.50 0.86 0.86 0.60 0.69 0.51
Support 0.25 0.55 0.06 0.73 0.63 0.30 0.60 0.33
Other 0.05 0.54 0.27 0.78 0.78 0.35 0.67 0.33
Weighted Average 0.45 0.60 0.20 0.83 0.81 0.48 0.67 0.43
VLMs struggle with nuanced disaster communication, underperforming compared to classic classifiers. Table 2 shows that VLMs, including LLaVA, GPT-4o-mini, and smolVLM, consistently underperform across most disaster-related classes. Their weighted F1-scores range from 0.20 (smolVLM) to 0.60 (GPT-4o-mini), all falling short of the best-performing baseline (Gaussian Naive Bayes, 0.67). Similarly, the Decision Tree classifier achieved a weighted score of 0.48, comparable to LLaVA’s 0.45. The performance of smolVLM in particular was notably poor, with a weighted average of only 0.20 and zero scores in multiple categories. These results suggest that, despite their general-purpose capabilities, current VLMs struggle to perform reliably in domain-specific, fine-grained classification tasks, such as the case for WildFireCan-MMD.
Custom classifiers are more robust and reliable. In contrast to VLMs, our custom classifiers (2-head and 3-head architectures) demonstrate superior and more stable performance. With weighted F1-scores of 0.83 and 0.81, they outperform all baselines and VLMs across nearly all categories. The 2-head classifier achieves top scores in Insurance (0.93), Reports of Actions of Responders (0.89), Smoke and Air Quality (0.86). The 3-head model benefits from incorporating OCR-extracted text, performing particularly well in Infrastructure (0.95), and Weather reports (0.91). While incorporating OCR-extracted textual information doesn’t improve the classification results overall, it might be a favorable approach when certain classes are deemed to be more critical for certain response strategies.
6 Analysis of Trends in Data
@@
After labeling our dataset and determining our best-performing classifier (the 2-head classifier described in section 4.3 and Figure D.1), we proceeded to collect another 46,279 posts from X. The collection methodology was the same as explained in section 3.1, where we gathered tweets using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. We combined this with a keyword-based search, then combined the results and removed duplicates. The full set of queries used for collecting this new data can be seen in Appendix F.
First, we extracted the author locations from the samples. Our goal was to obtain the province-level region where each sample is coming from. After extracting the author locations, which could be either automatically generated or user-provided, we ran the queries through OSMNX [^10] https://osmnx.readthedocs.io/en/stable/. This allowed us to automatically search for the locations of cities or towns that users provided and extract the province level information. If a users location was not found, or not in Canada, it was labeled as such. The distribution of posts by province can be seen in Figure 2, showing that the majority of data is posted from British Columbia and Alberta. Additional information about this data, including a word cloud of the text and the frequency distribution of the posts by word count can be found in Figures 1(a) and 1(b).
@@
Using our best-performing classifier with an F1-score of 83%, we obtained predictions for the new data. This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample. This of course, will not uphold the same gold standard of labeling that a human annotator could achieve, and manual labeling will have to be done in the future. Individual classifications should be approached with skepticism, but this method should still allow for general trends in the data to be analyzed on a broader scale.
@@
7 Discussion
This work benchmarks VLMs on a new dataset. A critical consideration in evaluation of the VLMs is that WildFireCan-MMD establishes a novel benchmark for evaluating these models. Most evaluation research works are performed with pre-existing datasets, and these datasets are also highly likely to be used for training VLMs and do not represent a genuine test set. Our newly created dataset evaluates the real generalization capabilities of such models since it was not available to the creators of these models during training. This is important for information management during crises because new disasters have unique requirements.
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. Also, smolVLM, while being both open-source and computationally affordable, has a performance far below acceptable rates. Therefore, when labeled data is unavailable, VLMs might be considered as a zero-shot classifier, but in practice, these models are not particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as decision tree and Gaussian NB, when trained with task-specific labeled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning common classification structures, such as pre-trained transformer-based text and image encoders. These findings align with other works such as Nadeem et al. (2025).
Custom classifiers can be used to discover social media use patterns. The primary function of the custom classifier is to perform real-time labeling of social media data during wildfire events, enabling the timely extraction of actionable information to support disaster response efforts. Beyond real-time deployment, our post-hoc analysis of a large unlabeled dataset illustrates the classifier’s capacity to identify latent patterns and temporal trends, providing a scalable tool for retrospective analysis and evidence-based policy development.
To enhance the accuracy and robustness of our model, future work can consider the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023). Fine-tuning of VLMs is also a promising avenue of improving the task which should be explored in future works Duan et al. (2024); Zhai et al. (2024).
@@ 8 Conclusion
Despite the advancements in general-purpose VLMs, they struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. This low accuracy, alongside the resource-intensive nature of such models, necessitates developing smaller fine-tuned models, specialized for specific regions and contexts. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs, and suitable for training specialized models. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
9 Acknowledgment
@@
While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
The first limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
@@
Appendix A Data collection Queries
@@
Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^11] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
After forming the clusters, BERTopic extracts keywords from each topic to uncover their central theme. For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.

### v3

@@ 4 Classification Methodology
This section describes a range of classifiers evaluated for the 13-class (12 categories + ‘Other’) classification of WildFireCan-MMD. We investigate three categories of classification methods. First, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers. This is the right choice when no training data is available. Then we examined classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) to investigate a case where labeled data is available but computation is limited. Finally, we experimented with a fine-tuned transformer-based classifier to assess its performance in the presence of labeled data and moderate computational resources. Further information on classical methods and VLMs can be found in Appendices E and F.
Architecture of custom classifier: To choose the architecture of the custom classifier, we experimented with the architectures proposed by various authors on the CrisisMMD dataset (presented in Appendix G) and evaluated them on the WildFireCAD-MMD dataset. These included CNNs and transformers for the image modality, as well as various BERT-based transformers for the text modality. After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in literature and included in Table G.1 were tested to improve the classification. More information about these experiments, as well as an ablation study, can be found in Appendix H.
The best performing classifier consisted of two parallel encoders, an image and text encoder. The image encoder is ViT333’google/vit-base-patch16-384’, and the text encoder is RoBERTa444’roberta-base’. The output CLS token of the last hidden state from each are sent through separate projection layers to reduce the dimensionality from 768 to 512. The output is passed to the transformer fusion module consisting of two encoder layers, each with d_model=512, nhead=8, dim_feedforward=2048, dropout=0.2. A final classification head projects the dimension from 512 to 256, which is followed by ReLU activation and dropout=0.2, then finally from 256 to num_classes. Figure I.1 shows the structure of this architecture.
Training and Evaluation: Three-fold cross validation with stratified sampling was used to train the model and report the results (mean $\pm$ std). The random seeds from splitting the data were kept for reproducibility. The hyperparameters used were a batch size of 8 and an image size of ($384\times{384}\times{3}$). Text sequences are padded to a max length of 144. Adam optimizer is used with a learning rate of $1e^{-5}$, weight_decay=0.01, and cross-entropy loss function. The model is fully trained after 10 epochs. The best model is saved based on f-weighted-average. The model was tested on the held-out test set (938 samples).
@@ 5 Experimental Results
Table 3 shows the calculated f-scores for the VLMs, custom classifier, and baseline models per class and as an average over all classes. Appendix J shows the confusion matrices of the custom classifiers and VLMs. All metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in the dataset.
Table 3: F1-scores per-class for zero-shot VLMs (preferred when no training data available), 3-fold cross validated f-scores for the custom architecture (when training data and moderate computation is available), and 5-fold cross validated f-scores for baseline classifiers (when training data and limited computation is available).
zero-shot VLMs Classifier Baseline Classifiers
Class LLaVA GPT-4o smolVLM Qwen DT GNB KNN SVM
Evacuees 17.65 37.21 25.84 12.70 88.55±3.20 49.42±6.66 44.44±4.65 44.39±4.50 38.42±3.18
General Information 19.05 20.37 00.00 21.88 82.04±3.54 32.49±4.63 38.05±8.41 32.88±6.67 39.35±10.31
Preparedness 57.14 62.50 22.58 67.69 66.02±4.10 48.04±6.96 62.55±5.53 55.69±5.91 54.51±2.33
Weather Reports 64.52 62.39 00.00 50.00 89.01±4.51 45.27±6.21 61.70±4.80 64.31±5.07 66.69±4.24
Warnings & Status Updates 34.94 48.00 00.00 42.73 93.19±2.02 43.35±1.27 45.87±3.35 59.32±2.68 56.99±2.14
Reports of Actions of Responders 26.09 69.09 00.00 63.89 81.10±2.14 49.43±2.23 51.37±4.50 54.35±6.08 61.72±5.73
Infrastructure 36.36 37.50 00.00 50.00 87.88±1.86 51.62±5.27 66.32±5.60 68.85±4.35 70.37±3.81
Political 69.09 77.86 00.00 85.25 83.73±0.29 51.58±5.78 66.59±4.43 55.02±5.48 60.53±4.27
Insurance 88.52 86.15 93.75 90.91 87.01±1.22 87.25±5.10 90.96±4.03 87.27±5.24 92.72±3.97
Advertisement 08.33 73.17 00.00 45.16 89.37±1.71 49.42±6.66 47.84±5.22 64.63±6.17 63.96±11.29
Smoke & Air Quality 72.46 75.38 45.48 65.96 84.20±6.11 57.79±3.33 71.38±1.20 74.93±1.17 76.24±2.51
Support 27.27 56.52 05.86 43.84 77.72±2.26 38.76±7.23 59.99±7.05 45.73±6.62 44.34±9.21
Other 05.22 57.58 25.36 52.46 81.29±1.00 40.44±4.48 53.68±5.21 50.58±2.45 56.33±2.27
Weighted Average 44.69 61.17 19.86 55.94 84.48±0.69 49.30±2.01 59.56±1.10 60.90±0.80 62.65±1.29
The custom classifier achieved the highest weighted average F1-score (84.48%), outperforming both zero-shot vision-language models and traditional baseline classifiers across all classes. From the results in Table 3 and Appendix J, the ‘Evacuees’ class had the lowest F1-score under the custom classifier (78.55%), likely due to high variability in both imagery and associated text. On the other hand, the ‘Insurance’ class achieved the highest performance (91.86%), reflecting the strong visual and textual consistency across samples. Other high-performing classes include ‘Political’ (89.50 ‘Smoke& Air Quality’ (89.35%), and ‘Advertisement’ (89.17%), which may be more visually or textually distinct. In contrast, mid-range scores are seen in classes like ‘Support’ (77.12%) and ‘Warnings & Status Updates’ (79.31%), which often contain dense textual overlays that may not align well with the image modality.
6 Discussion
Task-specific fine-tuning is critical: Table 3 shows that classical models fine-tuned on labeled data generally outperform zero-shot VLMs, with the custom model achieving  23% higher f-score than VLMs. Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix G.
VLMs involve tradeoffs: Zero-shot VLMs, while being a practical option when no labeled data is available, pose challenges in accuracy and cost.GPT-4o-mini offers strong results, but is costly and lacks transparency. Open-source models like LLaVA, Qwen, and smolVLM are more accessible but generally underperform, though Qwen is a notable exception, and trails GPT-4 by just 5%. Overall, VLMs may offer a solution when labeled data is unavailable, but their performance is comparable to traditional models such as SVM and KNN, which require training data but can be implemented with minimum computation.
Future Work: The dataset exhibits a significant class imbalance, and future work may involve steps to address this issue. Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research Sirbu et al. (2022); Sîrbu et al. (2025). Fine-tuning of VLMs is also a promising avenue which should be explored in future works Duan et al. (2024); Zhai et al. (2024).
7 Analysis on Trends in Data
@@
Here, we present an analysis of an additional 46,279 posts from X, posted between 2018 and 2024, using the custom classifier described in Section 4. The collection methodology was the same as explained in section 3.1, where tweets were gathered using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. This was combined with a keyword-based search, and then the data was combined, and duplicates were removed. The full set of search queries can be seen in Appendix K.
@@
The author locations were extracted from the samples with the goal of obtaining the province-level region from which each sample was posted. These could be either automatically generated or user-provided. The queries were run through OSMNX [^5] https://osmnx.readthedocs.io/en/stable/, which allowed for the automated search of cities or towns that users provided. If a user’s location was not found, or not in Canada, it was labeled as such. The distribution of posts by province is shown in Figure 2, indicating that the majority of the data originates from British Columbia and Alberta. Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix L.
Using the custom classifier, predictions were obtained for the new data. This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample. This of course, will not uphold the same gold standard of labeling that a human annotator could achieve.However, while individual classifications may not always be accurate, general trends in the data can be analyzed and linked to real events.
@@ 8 Conclusion
Despite the existence of other natural disaster datasets from social media, these datasets often lack wildfire data from a significant period of time, typically spanning only a single month surrounding a particular event. Furthermore, even less of this data comes from Canada and is mainly restricted to the US context. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation by multiple independent annotators. Unlike existing disaster datasets, WildFireCan-MMD is specifically designed for the Canadian wildfire context and provides a rich resource for training and evaluating specialized models. It also serves as a true test set for evaluating VLMs in a zero-shot setting. We provide the results of classifiers trained on this dataset from classical approaches to custom deep learning architectures and several VLMs, to showcase various scenarios in terms of data and computational resource availability. Our best results are obtained with an architecture that utilizes pre-trained encoders and transformer fusion for classification on this dataset, which highlights the importance of a labelled dataset. As climate-induced wildfires grow in frequency and complexity, this dataset and model support the development of robust tools for detection, coordination, and recovery.
Acknowledgment
@@
While this study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
The first limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of these results to other wildfire events or regions.
@@
Appendix A Data Collection
@@
Appendix B Annotation
Figure B.1 shows the distribution of the labels over the classes by each independent annotator.
Figure B.1: Counts of labels by each annotator
Appendix C Topic Modeling
Figure C.1 shows examples of topics extracted by the unsupervised BERTopic. We present representative images and corresponding keywords generated by BERTopic for each example. As seen in the captions of the image, some of the topics are not useful and present non-coherent concepts, while others are useful and relevant to wildfires.